OpenStudy (zubhanwc3):

a point is moving along the graph of y=1/(1+x^2)

OpenStudy (zubhanwc3):

a point is moving along the graph of y=1/(1+x^2) such that dx/dt = 2cm/sec(change in x in respect to time). find dy/dt for x=2 and x=-2

OpenStudy (tkhunny):

Have you considered the derivative? \(y = \dfrac{1}{1+x^{2}} = \left(1+x^{2}\right)^{-1}\) \(\dfrac{dy}{dt} = -\left(1+x^{2}\right)^{-2}\cdot 2x\cdot \dfrac{dx}{dt} = -\dfrac{2x}{1+x^{2}}\cdot\dfrac{dx}{dt}\)